OpenStudy (anonymous):

Solve the given equation 1269^2n=216^n+10

OpenStudy (anonymous):

very difficult, what topic is this?

OpenStudy (anonymous):

algebra 2, >< my worst subject. do you want answer choices?

OpenStudy (anonymous):

i meant it is quadratic equations? or exponectial functions?

OpenStudy (anonymous):

exponential function , i believe.

OpenStudy (anonymous):

the answer choices are a. 10/3 b. 2 c. 6 d. 10

OpenStudy (anonymous):

the equation is \[1269^{2n}=216^{n}+10?????\]

OpenStudy (anonymous):

>< no, n+10 is all in the superscript

OpenStudy (anonymous):

mmmmmmmmmmmm well i was solving a diferent problem then could u write the equation clearly?

OpenStudy (anonymous):

\[1269^{2n}=216^{n +10}\]

OpenStudy (anonymous):

If it helps, 1269=6^4 and 216 = 6^3

OpenStudy (anonymous):

yes it helps

OpenStudy (anonymous):

\[(6^{4})^{2n} = (6^{3})^{(n+10)}\]

OpenStudy (anonymous):

do u understand until there?

OpenStudy (anonymous):

Yes, o:

OpenStudy (anonymous):

i got an anwer diferent from ur choices let me check a minute

OpenStudy (anonymous):

mmkay, i'm not a math person. my forte is in language arts and science.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i check everysingle of ur possible answers and none of them are the solution

OpenStudy (anonymous):

i don't know, i'm going to guess then. -.-

OpenStudy (anonymous):

instead i check my answer and the equation become true

OpenStudy (anonymous):

hey buddy at least a thanks will be fine the ppl invest time helping u out and u dont even says tanks what education u got?